# SI-MIL: Taming Deep MIL for Self-Interpretability in Gigapixel Histopathology

### v1

@@ 10 SI-MIL ablation studies: components
We further investigate the necessity of deep features in SI-MIL. For this purpose, we substituted deep features with pathological features in the conventional MIL branch, thereby using the same pathological features in both SI-MIL branches. As shown in Table 4, the performance declines with or without $\mathcal{L}_{KD}$ when replacing deep features, underscoring the importance of employing potent deep features to guide the Self-Interpretable branch in SI-MIL.
@@ 11 Dataset contribution details (additional)
We contribute a unique comprehensive dataset aimed at enhancing interpretability and reproducibility in MIL research. The dataset encompasses nuclei maps, pathological features, and SI-MIL-generated patch-feature importance reports for over 2,200 WSIs. It covers multiple organs and cancer types, including Lung (lung adenocarcinoma vs. lung squamous cell carcinoma), Breast (invasive ductal vs. invasive lobular carcinoma), and Colon (low vs. high mutation). This diverse collection facilitates in-depth studies across various cancer types, providing a valuable resource for advancements in interpretable MIL methodologies.
Successful translation of AI tools to the clinic hinges upon the interpretability and trustworthiness of the tools. This dataset will serve as a critical asset for both the medical vision and digital pathology communities, facilitating the exploration of new research directions in the development of interpretable AI techniques for computational pathology. A significant obstacle in digital pathology research has been the intensive resource requirements for extracting features that possess clear geometric and physical significance, and which are interpretable by pathologists. The dataset creation involved analyzing gigapixel WSIs at 40$\times$ magnification, leveraging HoVer-Net [17] for cell segmentation and classification, followed by extracting pathological features and feature importance scores detailed in Sec. 3.3 and Sec. 5, respectively. Processing each WSI required $\sim$2 hours, divided between GPU-based cell map prediction and CPU-based pathological feature extraction. Employing three RTX 8000 GPUs and a 40-core CPU with 500GB RAM, the total processing amounted to $\sim$4400 hours ($\approx$60 days). We provide the comprehensive set to enable further research.
In view of $\sim$2 TB memory foorprint of HoVer-Net nuclei maps and the processed pathological features, we intend to host this dataset on TCIA Analysis Results, akin to other popular preprocessed datasets [25, 55].
@@ 13 Global interpretability analysis (addition)
Figure 9: Cohort-level Interpretation: Separability of top $K$ patches of WSIs across classes in the pathological feature space. Multivariate and Univariate analyses depict that the top $K$ patches selected by SI-MIL and their pathological features are more separable.
@@ 14 Top-$K$ comparative analysis
In this section, we demonstrate how the Self-Interpretable branch of SI-MIL tames the patch attention map of conventional deep MIL. Specifically, Figures 10 and 11 compare the spatial attention maps generated after the training of conventional MIL (i.e., without pathological features) and SI-MIL, which integrates both the conventional MIL and Self-Interpretable branches. We proceed to visualize the top $K=20$ patches from both MIL methods, categorizing them into groups based on whether they are common or distinct between the methods.
@@
15 Hand-crafted pathological feature extraction
In Sec. 3.3, we briefly discussed the categories of handcrafted pathological features, such as Morphometric and Spatial distribution properties. This section provides a detailed description of these features, accompanied by visualizations to elucidate the significance of their geometrical and physical meaning in computational pathology. We further refine these features into three categories, based on the studies from which they were adopted: Morphometric properties (from FLocK [35]), Graph-based Social Network Analysis [66], and Spatial heterogeneity properties [42].
@@ 15.2 Normalization
[…]

### v2

@@ 12 SI-MIL ablation studies: components
We further investigate the necessity of deep features in SI-MIL. For this purpose, we substituted deep features with PathExpert features in the conventional MIL branch, thereby using the same PathExpert features in both SI-MIL branches. As shown in Table 5, the performance declines with or without $\mathcal{L}_{KD}$ when replacing deep features, underscoring the importance of employing potent deep features to guide the Self-Interpretable branch in SI-MIL.
@@ 13 Dataset contribution details (additional)
We contribute a unique comprehensive dataset aimed at enhancing interpretability and reproducibility in MIL research. The dataset encompasses nuclei maps and PathExpert features for over 2,200 WSIs. SI-MIL-generated patch-feature importance reports will also be made available for representative slides. It covers multiple organs and cancer types, including Lung (lung adenocarcinoma vs. lung squamous cell carcinoma), Breast (invasive ductal vs. invasive lobular carcinoma), and Colon (low vs. high mutation). This diverse collection facilitates in-depth studies across various cancer types, providing a valuable resource for advancements in interpretable MIL methodologies.
Successful translation of AI tools to the clinic hinges upon the interpretability and trustworthiness of the tools. This dataset will serve as a critical asset for both the medical vision and digital pathology communities, facilitating the exploration of new research directions in the development of interpretable AI techniques for computational pathology. A significant obstacle in digital pathology research has been the intensive resource requirements for extracting features that possess clear geometric and physical significance, and which are interpretable by pathologists. The dataset creation involved analyzing gigapixel WSIs at 40$\times$ magnification, leveraging HoVer-Net [17] for cell segmentation and classification, followed by extracting PathExpert features and feature importance scores detailed in Sec. 3.2 and Sec. 5, respectively. Processing each WSI required $\sim$2 hours, divided between GPU-based cell map prediction and CPU-based PathExpert feature extraction. Employing three RTX 8000 GPUs and a 40-core CPU with 500GB RAM, the total processing amounted to $\sim$4400 hours ($\approx$60 days). We provide the comprehensive set to enable further research.
In view of $\sim$2 TB memory foorprint of HoVer-Net nuclei maps and the processed PathExpert features, we intend to host this dataset on TCIA Analysis Results, akin to other popular preprocessed datasets [25, 55].
@@ 15 Global interpretability analysis (addition)
Figure 9: Cohort-level Interpretation: Separability of top $K$ patches of WSIs across classes in the PathExpert feature space. Multivariate and Univariate analyses depict that the top $K$ patches selected by SI-MIL and their PathExpert features are more separable.
@@ 16 Top-$K$ comparative analysis
In this section, we demonstrate how the Self-Interpretable branch of SI-MIL tames the patch attention map of conventional deep MIL. Specifically, Figures 10 and 11 compare the spatial attention maps generated after the training of conventional MIL (i.e., without PathExpert features) and SI-MIL, which integrates both the conventional MIL and Self-Interpretable branches. We proceed to visualize the top $K=20$ patches from both MIL methods, categorizing them into groups based on whether they are common or distinct between the methods.
@@
17 Hand-crafted PathExpert feature extraction
In Sec. 3.2, we briefly discussed the categories of handcrafted PathExpert features, such as Morphometric and Spatial distribution properties. This section provides a detailed description of these features, accompanied by visualizations to elucidate the significance of their geometrical and physical meaning in computational pathology. We further refine these features into three categories, based on the studies from which they were adopted: Morphometric properties (from FLocK [35]), Graph-based Social Network Analysis [66], and Spatial heterogeneity properties [42].
@@ 17.2 Normalization
In this study, we implemented a two-step normalization process for all handcrafted PathExpert features. The first step addresses potential inaccuracies in nuclei segmentation and classification by HoVer-Net, using a binning operation. Each feature within a patch is assessed based on its percentile relative to other patches in the training split of a WSI task. These percentiles are then categorized into 10 discrete bins, ranging from the 0-10th to the 90-100th percentile, effectively shifting the scale of features from absolute values to a relative range from very-low to very-high. This approach transforms feature values into a robust and interpretable format across different patches. The second step involves mean and standard deviation normalization, once again using the training split data. This step centers the data around zero, optimizing it for effective processing by neural networks.